OpenStudy (anonymous):

f(x)=6x and g(x)=3x-7;f(x)/g(x)

TheSmartOne (thesmartone):

6x/3x-7

OpenStudy (anonymous):

(6x)/(3X-7)

TheSmartOne (thesmartone):

Good job @dannycobb5215

OpenStudy (anonymous):

its actually: (2-6x)/7

OpenStudy (anonymous):

Hi welcome to Openstudy! you are given: f(x)=6x g(x)=3x-7 and you are looking for \(\frac{f(x)}{g(x)}\), so: \(\LARGE \frac{f(x)}{g(x)}=\frac{6x}{3x-7}\)

TheSmartOne (thesmartone):

@dannycobb5215 it is not

OpenStudy (anonymous):

you cannot simplify this further

TheSmartOne (thesmartone):

@dannycobb5215 look at @Data_LG2 's explanation